How does the diaphragm work to change air pressure in the lungs, and as a result,...

  1. How does the diaphragm work to change air pressure in the lungs, and as a result, cause breathing?

  1. List the structures of the respiratory system in the order that they function for inhalation. Describe each structure’s function in detail.

  1. List the structures of the respiratory system in the order that they function for exhalation. Describe each structure’s function in detail.

  1. How is debris removed from the tracheobronchial tree?
  1. Define all pulmonary volumes and representative value for a healthy adult male. How do these values change for a healthy adult female?
  1. Define all pulmonary capacities and representative value for a healthy adult male. How do these values change for a healthy adult female?
  1. What is dead space? Contrast anatomical dead space with physiological dead space.
Answer #1

Diaphragmatic movement: During inspiration, as a result of discharge in phrenic neurons (c3, c4, c5), muscle fibres contract and draw the central tendon downwords by 1.5 cm inn eupnea; and by 7 cm in deep inspiration. This cause an increase in the vertical diameter of the thoracic cage. For each 1 cm descent 200-300 mL of air is sucked in, therefor diaphragmatic movements account for as much as 75% of tidal volume in eupnea.

MECHANIS OF INSPIRATION

Inspiration causes enlargement of thoracic cage. Thoracic cage enlarges because of increase in all diameters, viz. anteroposterior, transverse and vertical diameters. Anteroposterior and transverse diameters of thoracic cage are increased by the elevation of ribs. Vertical diameter is increased by the descent of diaphragm.

In general, change in the size of thoracic cavity occurs because of the movements of four units of structures:
1. Thoracic lid
2. Upper costal series
3. Lower costal series
4. Diaphragm.

1. Thoracic Lid
Thoracic lid is formed by manubrium sterni and the first pair of ribs. It is also called thoracic operculum. Movement of thoracic lid increases the anteroposterior diameter of thoracic cage. Due to the contraction of scalene muscles, the first ribs move upwards to a more horizontal position. This increases the anteroposterior diameter of upper thoracic cage.

2. Upper Costal Series
Upper costal series is constituted by second to sixth pair of ribs. Movement of upper costal series increases the anteroposterior and transverse diameter of the thoracic cage.
Movement of upper costal series is of two types:
i. Pump handle movement
ii. Bucket handle movement.

Pump handle movement: Contraction of external intercostal muscles causes elevation of these ribs and upward and forward movement of sternum. This movement is called pump handle movement. It increases anteroposterior diameter of
the thoracic cage.
Bucket handle movement: Simultaneously, the central portions of these ribs (arches of ribs) move upwards and outwards to a more horizontal position. This movement is called bucket handle movement and it increases the transverse diameter of
thoracic cage.


3. Lower Costal Series Lower costal series includes seventh to tenth pair of ribs. Movement of lower costal series increases the transverse diameter of thoracic cage by bucket handle movement.

Bucket handle movement: Lower costal series of ribs also show bucket handle movement by swinging outward and upward. This movement increases the transverse diameter of the thoracic cage.

Eleventh and twelfth pairs of ribs are the floating ribs. These ribs are not involved in changing the size of thoracic cage.

4. Diaphragm: SAME AS ABOVE

MECHANISM OF EXPIRATION

It is a passive process, but during forced expiration, the muscle of expiration contracts, which include

A. Anterior abdominal wall muscles

B. Internal intercoastal muscles

  1. contraction of anterior abdominal wall muscles increases intra-abdomina pressure and draws the lower ribs down and medially
  2. Internal intercostal muscles/; They pass obliquely downwards and posteriorly from rib to rib. on contraction they pull the upper ribs down so that the ribs aquire the position as seen at the end of expiration.

Debris removed from the tracheobronchial tree by the alveolar capillaries blood flow.

LUNG VOLUMES
Static lung volumes are the volumes of air breathed by an individual. Each of these volumes represents the volume of air present in the lung under a specified static condition (specific position of thorax). Static lung volumes are of four types:

  1. Tidal volume
  2. Inspiratory reserve volume
  3. Expiratory reserve volume
  4. Residual volume.

1.TIDAL VOLUME
Tidal volume (TV) is the volume of air breathed in and out of lungs in a single normal quiet respiration. Tidal volume signifies the normal depth of breathing.

Normal Value
500 mL (0.5 L).

2.INSPIRATORY RESERVE VOLUME
Inspiratory reserve volume (IRV) is an additional volume of air that can be inspired forcefully after the end of normal inspiration.
Normal Value
3,300 mL (3.3 L).

3.EXPIRATORY RESERVE VOLUME
Expiratory reserve volume (EVR) is the additional volume of air that can be expired out forcefully, after normal expiration.
Normal Value
1,000 mL (1 L).

4.RESIDUAL VOLUME
Residual volume (RV) is the volume of air remaining in lungs even after forced expiration. Normally, lungs cannot be emptied completely even by forceful expiration. Some quantity of air always remains in the lungs even after the forced expiration.

Residual volume is significant because of two reasons:
1. It helps to aerate the blood in between breathing and during expiration
2. It maintains the contour of the lungs.
Normal Value
1,200 mL (1.2 L)

(The volume of adult female lungs is typically 10-12% smaller than that of males who have the same height and age)

LUNG CAPACITIES
Static lung capacities are the combination of two or more lung volumes. Static lung capacities are of four types:
1. Inspiratory capacity
2. Vital capacity
3. Functional residual capacity
4. Total lung capacity.


1.INSPIRATORY CAPACITY
Inspiratory capacity (IC) is the maximum volume of air that is inspired after normal expiration (end expiratory position). It includes tidal volume and inspiratory reserve
volume (Fig. 121.1).
IC = TV + IRV
= 500 + 3,300 = 3,800 mL

2.VITAL CAPACITY (VC)
Vital capacity (VC) is the maximum volume of air that can be expelled out forcefully after a deep (maximal)
inspiration. VC includes inspiratory reserve volume, tidal volume and expiratory reserve volume.

VC = IRV + TV + ERV
= 3,300 + 500 + 1,000 = 4,800 mL
Vital capacity is significant physiologically and its determination is useful in clinical diagnosis as explained later in this chapter.

3.FUNCTIONAL RESIDUAL CAPACITY
Functional residual capacity (FRC) is the volume of air remaining in lungs after normal expiration (after normal tidal expiration). Functional residual capacity includes expiratory reserve volume and residual volume.

FRC = ERV + RV
= 1,000 + 1,200 = 2,200 mL

4.TOTAL LUNG CAPACITY
Total lung capacity (TLC) is the volume of air present in lungs after a deep (maximal) inspiration. It includes all the volumes.
TLC = IRV + TV + ERV + RV
= 3,300 + 500 + 1,000 + 1,200 = 6,000 mL

(The volume and capacities of adult female lungs is typically 10-12% smaller than that of males who have the same height and age)

DEAD SPACE

Dead space is defined as the part of the respiratory tract, where gaseous exchange does not take place. Air present in the dead space is called dead space air.

Dead space is of two types:
1. Anatomical dead space
2. Physiological dead space.

Anatomical Dead Space
Anatomical dead space extends from nose up to terminal bronchiole. It includes nose, pharynx, trachea,
bronchi and branches of bronchi up to terminal bronchioles. These structures serve only as the
passage for air movement. Gaseous exchange does not take place in these structures.

Physiological Dead Space
Physiological dead space includes anatomical dead space plus two additional volumes.

Additional volumes included in physiological dead space are:
1. Air in the alveoli, which are non-functioning. In some respiratory diseases, alveoli do not function because of dysfunction or destruction of alveolar membrane.
2. Air in the alveoli, which do not receive adequate blood flow. Gaseous exchange does not take place
during inadequate blood supply